OpenStudy (anonymous):

76 is 40% of (what number)______?

OpenStudy (anonymous):

Use the same rule of = times % = x/100 or in this case 40/100

OpenStudy (anonymous):

and is ---> =

OpenStudy (anonymous):

0.4?

OpenStudy (anonymous):

40/100 = 0.4, yes

OpenStudy (anonymous):

But to find the answer, you need to use 40/100 times x = 76

OpenStudy (anonymous):

which gives you 190

OpenStudy (anonymous):

\[{40 \over 100} \times x = 76\]\[40x=76 \times 100\]\[x={7600 \over 40}\]\[=190\]